# A Categorical Framework for Quantifying Emergent Effects in Network Topology

### v3

@@ Abstract
Emergent effect is crucial to understanding the properties of complex systems that do not appear in their basic units, but there has been a lack of theories to measure and understand its mechanisms. In this paper, we established a framework based on homological algebra that encodes emergence as the mathematical structure of cohomologies and then applied it to network models to develop a computational measure of emergence. This framework ties the emergence of a system to its network topology and local structures, paving the way to predict and understand the cause of emergent effects. We show in our numerical experiment that our measure of emergence correlates with the existing information-theoretic measure of emergence.
@@ 1 Introduction
The paper will be organized as follows: In Section 2 we define and illustrate emergent effects through an example. In Section 3, we present the framework for emergence as loss of exactness, adapted from Adam (2017). In Section 4 and 5, this framework is realized as the mapping between networks, or more generally, quiver representations and then we compute the mathematical structure which encodes emergence. In Section 6 we establish the numerical measure of emergence and then discuss the numerical result showing a correlation with existing measures of emergence.
2 A Descriptive Framework for Emergent Effects
Two key conceptual components are necessary to qualitatively describe emergent effects within the framework proposed by Adam (see Adam, 2017). The first is a notion of interaction or interconnection between the components of a system. For example, a thermodynamics system consisting of particles exhibits interesting properties because of the interactions between particles. The second is the notion of interactional effects, which equips each system with an observable. For example, we perceive and can feel hot or cold, and can measure temperature, due to the interactions between particles, but not from the observation or description of any number of individual particles. These kinds of interactional effects are almost always associated with partial observations, or a simplification and integration of lower — more foundational or granular —levels or scales in the system that result in a ’loss of information’ or inability to properly observe the effects of those interactions at a higher level.
@@ 2 A Descriptive Framework for Emergent Effects
To formalize these ideas, we begin by representing the interactions between components as an operation $\vee$, where $s_{1}\vee s_{2}$ represents a new interconnected system of subsystems $s_{1}$ and $s_{2}$. Interactional effects are described by the mapping $\Phi$ that sends a system to its partial observation or interactional effect at a higher scale, in some cases corresponding to a coarse graining scheme rosas2024software (Rosas et al., 2024). Emergent effects are sustained whenever the observation of the combined system cannot be explained by the observation of the separate components. Mathematically then,
@@
We will discuss how to measure the difference between $\Phi(s_{1}\vee s_{2})$ and $\Phi(s_{1})\vee\Phi(s_{2})$ in Section 2. We hereby provide three examples, in machine learning, biology and statistical mechanics to illustrate this idea of emergence.
@@ 2.1 Machine learning
Emergence or generativity has been a rising concept in machine learning (for example, Wei et al., 2022; wei2022chain, Wei et al., 2022, Du et al., 2024). Emergent abilities of large language models, for example, Wei et al. (2022), commonly conceived as the new properties/ abilities of the larger models that do not exist in smaller models. If we consider $s_{1}$ and $s_{2}$ as two smaller models, $s_{1}\vee s_{2}$ as combining two smaller models into a larger model by, for example, techniques in ensemble learning, (see, e.g., Mohammed & Kora, 2023), and $\Phi$ as the mapping that reflects the properties/ abilities of the model, that is, $\Phi(s)$ is the ability acquired by the model $s$. Then $\Phi(s_{1}\vee s_{2})$ is the properties/ abilities of the combined model and $\Phi(s_{1})\vee\Phi(s_{2})$ can be interpreted as a summation of the properties/ abilities of each small model. Then the difference between $\Phi(s_{1}\vee s_{2})$ and $\Phi(s_{1})\vee\Phi(s_{2})$ can reflect the emergent properties/ abilities that result in the nonlinear increase of performance, related to the performance in Wei et al. (2022). The difference can also be related to generalizability, where $s_{1}$ and $s_{2}$ are two data sets, when the model trains on two data sets, it is usually different from training the model on separate datasets.
@@ 2.2 Biology
In biological systems, the concept of emergence is exemplified through the self-organizing behaviors observed in social insects, particularly in ant colonies. Individual ants, operating via a set of simple rules driven by genetic programming and environmental interaction, primarily pheromone communication, exhibit relatively basic behaviors when isolated. However, when these individuals interact as part of a larger group, the collective behavior of the colony transcends the capabilities of any single ant. The emergent properties of the colony, such as complex nest architecture, efficient foraging strategies, and intricate social organization, arise from the dynamic interplay and feedback loops among the ants. Each ant’s contributions, though simple in isolation, collectively result in a higher-level organization that is qualitatively different from and cannot be predicted merely by understanding the behavior of individual ants. Consider $s_{1}$ and $s_{2}$ as two entities, for example, two ant groups, and $s_{1}\vee s_{2}$ represents the large group formed by $s_{1}$ and $s_{2}$. Now let $\Phi(s)$ represent the colony structure formed by ant group $s$. Then $\Phi(s_{1}\vee s_{2})$ represents the colony structure built through the interaction among the ants, which exhibits complex structures compared to $\Phi(s_{1})\vee\Phi(s_{2})$, the colony structure built without the interaction between $s_{1}$ and $s_{2}$. The difference between $\Phi(s_{1}\vee s_{2})$ and $\Phi(s_{1})\vee\Phi(s_{2})$ reflects the emergence exhibited in the example of ant colonies.
@@ 2.3 Statistical Mechanics
In the domain of statistical physics, the concept of emergent phenomena is profoundly illustrated by the study of phase transitions, a field where macroscopic changes emerge from the collective dynamics of microscopic constituents. Phase transitions, such as the crystallization of a liquid or the magnetization of a ferromagnetic material, represent a paradigmatic example of how large-scale, qualitative changes in a system’s properties can arise from the cooperative interactions among its many individual components. Theoretical models, such as the Ising model for ferromagnetism, exemplify this: individual atomic spins interact locally, but when the system reaches a critical temperature, a dramatic, collective alignment occurs, manifesting as a spontaneous magnetization. Similarly, in the process of freezing, the molecular disorder of a liquid yields to an ordered crystalline structure as temperature decreases. These transitions are governed by principles like symmetry breaking and the minimization of free energy. For example, consider the 2D Ising system, phase transition happens at the discontinuity of $M$ the total magnitization of the system as the external field $J$ varies. We can represent $s_{1}$ as the Ising system and $s_{2}$ as the system that produced the external field, and $s_{1}\vee s_{2}$ as the coupled system. We can represent $\Phi(s_{1})$ as the total magnetization of the system $s_{1}$. Now we have that $\Phi(s_{1}\vee s_{2})$ represents the total magnetization of the coupled system, and $\Phi(s_{1})\vee\Phi(s_{2})$ represents the average total magnetization of individual systems. The difference between $\Phi(s_{1}\vee s_{2})$ and $\Phi(s_{1})\vee\Phi(s_{2})$ represents the susceptibility of the system under the change of external fields which captures phase transition.
@@ 3 Emergence as loss of exactness
In this section, we reformulate the intuitive concepts of emergence introduced in the previous sections to lead to an effective quantitative measure. It turns out to be necessary to introduce some concepts of category theory and homological algebra, but let us show the intuitive idea in the beginning.
We say that a system is emergent if there exist two components satisfying the condition that $\Phi(s_{1}\vee s_{2})\neq\Phi(s_{1})\vee\Phi(s_{2})$. However, as discussed in the Introduction, on its own this definition of emergence is not sufficient to evaluate the amount of emergent effects that can present in a system. We need to approximate the difference when $\Phi(s_{1}\vee s_{2})$ and $\Phi(s_{1})\vee\Phi(s_{2})$ are representing two non-numerical structures, for example, networks. We show in Section 5 that a mapping $\Phi$ that sends a network to its partial structure can be realized as a functor, a basic concept in category theory.
Let’s consider the simple case when $\Phi$ is simply a smooth function $f:\mathbb{R}\rightarrow\mathbb{R}$ and the interaction $\vee$ is simply taking the average, $s_{1}\vee s_{2}=(s_{1}+s_{2})/2$. Then we realize that the extent to each $\Phi(s_{1}\vee s_{2})$ differs from $\Phi(s_{1})\vee\Phi(s_{2})$ is just $\Big{|}f(\frac{s_{1}+s_{2}}{2})-\frac{f(s_{1})+f(s_{2})}{2}\Big{|}$, which is related to how nonlinear the function $f$ is, and can be studied by the derivatives of $f$, in particular, the second order derivative, since $\Big{|}f(\frac{s_{1}+s_{2}}{2})-\frac{f(s_{1})+f(s_{2})}{2}\Big{|}$ can be approximated by $\frac{|s_{2}-s_{1}|^{2}}{4}|f^{\prime\prime}(\xi)|$ for some $\xi$ between $s_{1}$ and $s_{2}$. Now when $\Phi$ is a functor, which captures the cross-scale information flow in real world systems, we want an analogue to derivatives to apply this idea, and this naturally leads to the concept of a derived functor in homological algebra. We can also see that Definition 2.1 captures the structural nonlinearity in emergence, the nonlinearity of system’s behavior and functionality as the system’s structure changes, or as we go from components to parts of the system to the whole system.
@@
A category is an algebraic structure consisting of objects and morphisms. We consider the collection of subsystems and their constituent components that give rise to emergent effects as the category System, where each object is a subsystem. The morphisms represent a relation between subsystems to be interpreted, for example, $f:V\rightarrow W$ can give a way to embed subsystem $V$ into $W$. When subsystems interact they form a new subsystem (we also call it interconnection). This new resultant subsystem is captured as the mathematical concept of colimit with respect to the interaction diagram, which is treated in detail in Fong & Spivak (2018). Colimit is an object in System, as defined below, that satisfies certain properties. In Section 4 we discuss the category of quiver representations as System and we will give the colimit and see how it can be regarded as interconnection among objects.
@@
Within this framework of homological algebra, the loss of exactness can be encoded in the derived functor introduced below. This approach evaluates emergence via the algebraic properties of the functor itself. Given a function (or more generally a functor) $\Phi$ on a set (category) of networks that represents the processing of information in a system, we can extract information from the ”derivative” of the function (derived functor), represented as $L^{1}\Phi$ or $R^{1}\Phi$ in the standard literature. Now we show they encode information related to emergence.
@@
then $R^{i}\Phi(A)$ is defined as the $i$th cohomology object of the complex (similarly, $L^{i}\Phi(A)$ is defined as the $i$th cohomology object of the complex recovered from the projective resolution). Standard textbooks in homological albgera, for example, Rotman & Rotman (2009) has shown that they are indeed functors, and under the assumption that $\Phi$ is left-exact, we have $R^{0}\Phi(A)=\Phi(A)$. In Section 5 we will give a concrete definition and example that finds $L^{1}\Phi(A)$ and $R^{1}\Phi(A)$. Now note that the chain complex (8) we recovered is not necessarily exact. To measure how much exactness is lost under this functor, we can take advantage of the following standard result in homological algebra (see, e.g. Gelfand & Manin, 2013), which gives us a long exact sequence
@@
To wrap up, in this section we introduced the homological algebra framework to measure emergence through derived functor, which is an analogue of derivatives in calculus. We note that derived functor can indeed be a generalization of derivatives: derived functor measures the extent to which a functor fails to preserve interactional effect (realized as colimit), analogous to how second-order or higher-order derivatives measure the extent to which a function fails to be linear. The nonlinearty in the functor’s behavior, encoded in the derived functor $R^{1}\Phi(A)$, reflects the nonlinearity in the system’s behavior as we go from single components to parts of the system to system itself, hence captures the concept of emergence.
@@ 4 Quiver representations
Quiver representation has the following advantages: first, the category of quiver representations $\textbf{Rep}(Q)$ is itself an Abelian category, the kind of category on which we can carry out homological algebra constructions. Second, plenty of mathematical theories on quiver representations and quiver varieties (see, e.g., Kirillov Jr, 2016; Jeffreys & Lau, 2022) have been developed that can potentially be applied to the formulation of neural networks. Third, the representation of nodes and edges can potentially be used to model information flow on the network and encode the biophysics that govern those networks.
Formally, a quiver is a directed graph where loops and multiple arrows between two vertices are allowed, defined as follows:
@@
We get a representation $V$ of $Q=(Q_{0},Q_{1},h,t)$ if we attach to every vertex $x\in Q_{0}$ a finite dimentional $\mathbb{C}$-vector space $V(x)$ and to every arrow $a\in Q_{1}$ a $\mathbb{C}$-linear map $V(a):V(ta)\rightarrow V(ha)$.
@@
We now want to formulate a collection of quiver representations into a category in order to be able to apply a homological algebra framework. We define a category ${\rm Rep}(Q)$ where the objects are representations of $Q$. For any two representations $V$ and $W$ $Q=(Q_{0},Q_{1},h,t)$. We define a morphism $\phi:V\rightarrow W$ by attaching to every vertex $x\in Q_{0}$ a $\mathbb{C}$ linear map $\phi(x):V(x)\rightarrow W(x)$ such that for every $a\in Q_{1}$ the diagram
@@
This result is proven as Proposition 1.5.4 in Derksen & Weyman (2017). It implies that every quiver representation, including network structures, can be interpreted as a module, an algebraic structure. The category of quiver representations, therefore, behaves similarly as the category of modules. Hence, $\textbf{Rep}(Q)$ is indeed an Abelian category, (see, Derksen & Weyman, 2017), so the constructions of homological algebra are available. Thus, we can use tools primarily from linear algebra and graph theory, to compute emergence as the algebraic structure arising from the interactions between quiver representations.

### v4

@@ Abstract
Emergent effect is crucial to understanding the properties of complex systems that do not appear in their basic units, but there has been a lack of theories to measure and understand its mechanisms. In this paper, we consider emergence as a kind of structural nonlinearity, discuss a framework based on homological algebra that encodes emergence as the mathematical structure of cohomologies, and then apply it to network models to develop a computational measure of emergence. This framework ties the potential for emergent effects of a system to its network topology and local structures, paving the way to predict and understand the cause of emergent effects. We show in our numerical experiment that our measure of emergence correlates with the existing information-theoretic measure of emergence.
@@ 1 Introduction
The paper will be organized as follows: In Section 2 we define and illustrate emergent effects through examples. In Section 3, we present the framework for emergence as loss of exactness, adapted from Adam (2017). In Section 4 and 5, this framework is realized as the mapping between networks, or more generally, quiver representations and then we compute the mathematical structure which encodes emergence. In Section 6 we establish the numerical measure of emergence and then discuss the numerical result showing a correlation with existing measures of emergence.
2 Emergence as Structural Nonlinearity
Two key conceptual components are necessary to qualitatively describe emergent effects within the framework proposed by Adam (see Adam, 2017). The first is a notion of interaction or interconnection among the components of a system. For example, a thermodynamics system consisting of particles exhibits interesting properties because of the interactions among particles. The second is the notion of interactional effects, which equips each system with an observable. For example, we can measure global properties of particles systems, like temperature, but not from the observation or description at the level of individual particles. These kinds of interactional effects are almost always associated with partial observations, or a simplification and integration of lower — more foundational or granular —levels or scales in the system that result in a ’loss of information’ or inability to properly observe the effects of those interactions at a higher level.
@@ 2 Emergence as Structural Nonlinearity
To formalize these ideas, we begin by representing the interactions between components as an operation $\vee$, where $s_{1}\vee s_{2}$ represents a new interconnected system of subsystems $s_{1}$ and $s_{2}$. Interactional effects are described by the mapping $\Phi$ that sends a system to its partial observation or interactional effect at a larger scale, in some cases corresponding to a coarse-graining scheme Rosas et al. (2024). Emergent effects are sustained whenever the observation of the combined system cannot be explained by the observation of the separate components. Mathematically,
@@
Let’s consider the simple case when $\Phi$ is simply a smooth function $f:\mathbb{R}\rightarrow\mathbb{R}$ and the interaction $\vee$ is simply taking the average, $s_{1}\vee s_{2}=(s_{1}+s_{2})/2$. Then we realize that the extent to which $\Phi(s_{1}\vee s_{2})$ differs from $\Phi(s_{1})\vee\Phi(s_{2})$ is just $\Big{|}f(\frac{s_{1}+s_{2}}{2})-\frac{f(s_{1})+f(s_{2})}{2}\Big{|}$, which is related to how nonlinear the function $f$ is, and can be studied by the derivatives of $f$, in particular, the second order derivative, since $\Big{|}f(\frac{s_{1}+s_{2}}{2})-\frac{f(s_{1})+f(s_{2})}{2}\Big{|}$ can be approximated by $\frac{|s_{2}-s_{1}|^{2}}{4}|f^{\prime\prime}(\xi)|$ for some $\xi$ between $s_{1}$ and $s_{2}$. Now when $\Phi$ models the cross-scale information flow in real world systems, we want an analogue to derivatives to apply this idea, and this naturally leads to the concept of a derived functor in homological algebra, which will be discussed in the next Section. We can also see that Definition 2.1 captures the structural nonlinearity in emergence, the nonlinearity of system’s behavior and functionality as the system’s structure changes, or as we go from components to parts of the system to the whole system.
@@
We will discuss the theoretical framework for measuring the difference between $\Phi(s_{1}\vee s_{2})$ and $\Phi(s_{1})\vee\Phi(s_{2})$ in Section 2, and the final measure will be given in Section 6. We hereby provide three examples, in machine learning, biology and statistical mechanics to illustrate this idea of emergence.
@@ 2.1 Machine Learning
Emergence, or generativity, is an increasingly important concept in machine learning (for example, Wei et al., 2022, 2022; Du et al., 2024). Emergent abilities of large language models, for example, Wei et al. (2022), commonly conceived as the new properties/ abilities of the larger models that do not exist in smaller models. If we consider $s_{1}$ and $s_{2}$ as two smaller models, $s_{1}\vee s_{2}$ as combining two smaller models into a larger model by, for example, techniques in ensemble learning, (see, e.g., Mohammed & Kora, 2023), and $\Phi$ as the mapping that reflects the properties/ abilities of the model, that is, $\Phi(s)$ is the ability acquired by the model $s$. Then $\Phi(s_{1}\vee s_{2})$ is the properties/ abilities of the combined model and $\Phi(s_{1})\vee\Phi(s_{2})$ can be interpreted as a summation of the properties/ abilities of each small model. Then the difference between $\Phi(s_{1}\vee s_{2})$ and $\Phi(s_{1})\vee\Phi(s_{2})$ can reflect the emergent properties/ abilities that result in the nonlinear increase of performance, related to the performance in Wei et al. (2022). The difference can also be related to generalizability, where $s_{1}$ and $s_{2}$ are two data sets, when the model trains on two data sets, it is usually different from training the model on separate datasets.
@@ 2.2 Biology
In biological systems, the concept of emergence is exemplified through the self-organizing behaviors observed in social insects, such as ants or bees Moffett et al. (2021). The behaviors of individual ants are governed by a set of simple rules driven by genetic programming and environmental interactions. Pheromones in particular are key chemicals involved in communication and the exchange of information between ants in a colony. However, when these individuals interact as part of a larger group, the collective behavior of the colony transcends the capabilities of any single ant. The emergent properties of the colony, such as complex nest building, efficient foraging strategies, and intricate social organization, arise from the dynamic interplay and feedback loops between the ants. Each ant’s contributions, though simple in isolation, collectively result in a higher-level organization that is qualitatively different from and cannot be predicted merely by understanding the behavior of individual ants. Consider $s_{1}$ and $s_{2}$ as two entities, for example, two groups of ants, and $s_{1}\vee s_{2}$ represents the large group formed by $s_{1}$ and $s_{2}$. Let $\Phi(s)$ represent the colony structure formed by the ant group $s$. Then $\Phi(s_{1}\vee s_{2})$ represents the colony structure built through the interaction among the ants, which exhibits complex structures compared to $\Phi(s_{1})\vee\Phi(s_{2})$, the colony structure built without the interaction between $s_{1}$ and $s_{2}$. The difference between $\Phi(s_{1}\vee s_{2})$ and $\Phi(s_{1})\vee\Phi(s_{2})$ reflects the emergence exhibited in the example of ant colonies.
@@ 2.3 Statistical Mechanics
In the domain of statistical physics, the concept of emergent phenomena is profoundly illustrated by the study of phase transitions Goldenfeld (2018), where macroscopic changes emerge from the collective dynamics of microscopic constituents. Phase transitions, such as the crystallization of a liquid or the magnetization of a ferromagnetic material, represent a paradigmatic example of how large-scale, qualitative changes in a system’s properties can arise from the cooperative interactions among its many individual components. Theoretical models, such as the Ising model for ferromagnetism, exemplify this: individual atomic spins interact locally, but when the system reaches a critical temperature, a dramatic, collective alignment occurs, manifesting as a spontaneous magnetization. Similarly, in the process of freezing, the molecular disorder of a liquid yields to an ordered crystalline structure as temperature decreases. These transitions are governed by principles like symmetry breaking and the minimization of free energy Goldenfeld (2018). For example, consider the Ising system, phase transition happens at the discontinuity of $M$ the total magnitization of the system as the external field $J$ varies. We can represent $s_{1}$ as the Ising system and $s_{2}$ as the system that produced the external field, and $s_{1}\vee s_{2}$ as the coupled system. We can represent $\Phi(s_{1})$ as the total magnetization of the system $s_{1}$. Now we have that $\Phi(s_{1}\vee s_{2})$ represents the total magnetization of the coupled system, and $\Phi(s_{1})\vee\Phi(s_{2})$ represents the average total magnetization of individual systems. The difference between $\Phi(s_{1}\vee s_{2})$ and $\Phi(s_{1})\vee\Phi(s_{2})$ represents the susceptibility of the system under the change of external fields which captures phase transition.
@@ 3 Emergence as Loss of Exactness
In this section, we reformulate the intuitive concepts of emergence introduced in the previous sections to derive a quantitative measure of emergence. The technical arguments necessitate an introduction to concepts of category theory and homological algebra. However, we begin by presenting the intuitive ideas.
In Definition 2.1 we say that a system is emergent if there exist two components satisfying the condition that $\Phi(s_{1}\vee s_{2})\neq\Phi(s_{1})\vee\Phi(s_{2})$. However, as discussed in the Introduction, on its own this definition of emergence is not sufficient to evaluate the amount of emergent effects that can present in a system. We need to approximate the difference when $\Phi(s_{1}\vee s_{2})$ and $\Phi(s_{1})\vee\Phi(s_{2})$ are representing two non-numerical structures, for example, networks. We show in Section 5 that a mapping $\Phi$ that sends a network to its partial structure can be realized as functors, a fundamental concept in category theory. We can use derived functors to capture the nonlinearity that exists in such functors.
@@
A category is an algebraic structure consisting of objects and morphisms. We consider the collection of subsystems and their constituent components that give rise to emergent effects as the category System, where each object is a subsystem. The morphisms represent a relation between subsystems to be interpreted. For example, $f:V\rightarrow W$ can give a way to embed subsystem $V$ into $W$. When subsystems interact they form a new subsystem (we also call it interconnection). This new resultant subsystem is captured as the mathematical concept of colimit with respect to the interaction diagram, which is treated in detail in Fong & Spivak (2018). Colimit is an object in System, as defined below, that satisfies certain properties. In Section 4 we discuss the category of quiver representations as System and we will give the colimit and see how it can be regarded as interconnection among objects.
@@
Within this framework of homological algebra, the loss of exactness can be encoded in the derived functor introduced below. This approach evaluates emergence via the algebraic properties of the functor itself. Given a function (or more generally a functor) $\Phi$ on a set (category) of networks that represents the processing of information in a system, we can extract information from the ”derivative” of the function (derived functor), represented as $L^{1}\Phi$ or $R^{1}\Phi$ in the standard literature. We next show that they encode information related to emergence.
@@
then $R^{i}\Phi(A)$ is defined as the $i$th cohomology object of the complex (similarly, $L^{i}\Phi(A)$ is defined as the $i$th cohomology object of the complex recovered from the projective resolution). Standard textbooks in homological albgera, for example, Rotman & Rotman (2009) have shown that they are indeed functors, and under the assumption that $\Phi$ is left-exact, we have $R^{0}\Phi(A)=\Phi(A)$. In Section 5 we will give a concrete definition and example that finds $L^{1}\Phi(A)$ and $R^{1}\Phi(A)$. Now note that the chain complex (8) we recovered is not necessarily exact. To measure how much exactness is lost under this functor, we can take advantage of the following standard result in homological algebra (see, e.g. Gelfand & Manin, 2013), which gives us a long exact sequence,
@@
In this section we introduced the homological algebra framework to measure emergence through derived functor, which is an analogue of derivatives in calculus. We note that derived functor can indeed be a generalization of derivatives: derived functor measures the extent to which a functor fails to preserve interactional effect (realized as colimit), analogous to how second-order or higher-order derivatives measure the extent to which a function fails to be linear. The nonlinearty in the functor’s behavior, encoded in the derived functor $R^{1}\Phi(A)$, reflects the nonlinearity in the system’s behavior as we go from single components to parts of the system to system itself, hence captures the concept of emergence.
@@ 4 Quiver Representations
The use of quiver representations as the System category has the following advantages: first, the category of quiver representations $\textbf{Rep}(Q)$ is itself an Abelian category, the kind of category on which we can carry out homological algebra constructions. Second, plenty of mathematical theories on quiver representations and quiver varieties have been developed that can potentially be applied to the formulation of neural networks (see, e.g., Kirillov Jr, 2016; Jeffreys & Lau, 2022). Third, the representation of nodes and edges can potentially be used to model information flow on the network and encode the biophysics that govern those networks.
A quiver is a directed graph where loops and multiple arrows between two vertices are allowed, and the representation of a quiver associates nodes with vector spaces and arrows with linear maps (see Figure 1, 2). Formally, they are defined as follows:
@@
We get a representation $V$ of $Q=(Q_{0},Q_{1},h,t)$ if we attach to every vertex $x\in Q_{0}$ a finite dimensional $\mathbb{C}$-vector space $V(x)$ and to every arrow $a\in Q_{1}$ a $\mathbb{C}$-linear map $V(a):V(ta)\rightarrow V(ha)$.
@@
We now want to formulate the collection of quiver representations into a category in order to apply the homological algebra tools. The category $\textbf{Rep}(Q)$ is defined where the objects are representations of $Q$. For any two representations $V$ and $W$ which are representations of quiver $Q=(Q_{0},Q_{1},h,t)$, we define a morphism $\phi:V\rightarrow W$ by attaching to every vertex $x\in Q_{0}$ a $\mathbb{C}$ linear map $\phi(x):V(x)\rightarrow W(x)$ such that for every $a\in Q_{1}$ the diagram
@@
This result is proven as Proposition 1.5.4 in Derksen & Weyman (2017). It implies that every quiver representation, including network structures, can be interpreted as a module, an algebraic structure. The category of quiver representations, therefore, behaves similarly as the category of modules. Hence, $\textbf{Rep}(Q)$ is indeed an Abelian category, (see, Derksen & Weyman, 2017), so the constructions of homological algebra are available. We can then use tools primarily from linear algebra and graph theory, to compute emergence as the algebraic structure arising from the interactions between quiver representations.